# Dynamic read & write optimizaton with TurtleKV

### v1

@@ 3.3. Checkpoint Distance
The $\chi$ parameter controls how frequently newly created pages (nodes, segments, and leaves) are written from memory to storage. At $\chi=1$, every time we perform a batch update on the tree, all modified pages must be written back out. At $\chi=2$, we defer the writing of new/modified pages until two batches have been applied, etc. It may not be immediately obvious to the reader that $\chi$ should have any effect on the write cost of key updates. To see that it does, let us revisit the batch insertion example illustrated by Figure 5, considering the span of time (as measured by batch number) that each page is live, shown in Figure 6. The pages which appear in durable storage at a given time step are those whose lifespan overlaps a given column in the middle section (Lifetime) of the diagram.
3.4. Cost Analysis & Limitations
Table 1 compares TurtleTrees to other data structures. We see that TurtleTree’s update put (update) operations benefit greatly from checkpoint distance due to the $\chi$ parameter in the denominator. Gets under different memory models perform similarly to other data structures, but are greatly increased by read caching (not shown in the table). We next describe how the table values are derived.
Individual key updates travel from root to leaf of the tree as batch updates and flushes take place. Since there are $O\left(\log_{2}{\rho}\right)$ levels within the buffer of each node, and the tree height is $O\left(\log_{\rho}{\left(N/L\right)}\right)$, each key update is written $O\left(\log_{2}{\rho}\cdot\log_{\rho}{\left(N/L\right)}\right)=O\left(log_{2}{\left(N/L\right)}\right)$ times, regardless of the node page size or tree branching factor $\rho$. Note that in order to correctly analyze the cost of a key update, we need only consider its initial journey from root-to-leaf, since each time a key’s page is re-written after it reaches the bottom of the tree, we can amortize that cost against the incoming batch, which represents an equal number/size of keys ending their root-to-leaf journey. If we define the transfer size $B$ in units equal to a single key update, the I/O cost of a key update is
$O\left(\frac{1}{B}\cdot\log_{2}{\frac{N}{L}}\right)$
Point queries (Get) differ in their I/O cost depending on whether or not the key is found. Here we give a worst-case bound for the general case; in practice this can be dramatically improved through the addition of per-leaf filters. The cost analysis here also depends on the specific I/O model used; if we are counting total page loads, as in the Disk Access Model (DAM), then a single-key (point) query must load all node pages and buffer segments along a single root-to-leaf path: $O\left(\log_{\rho}{\frac{N}{L}}+\frac{L}{B}\cdot\log_{2}{\frac{N}{L}}\right)$ or just $O\left(\frac{L}{B}\cdot\log_{2}{\frac{N}{L}}\right)$.
However, because “latency lags bandwidth” (Patterson, 2004), it is also reasonable to consider I/O models where latency and parallelism are explicitly accounted for (Bender et al., 2021, 2019; Aggarwal and Vitter, 1988), e.g. the Parallel Disk Access Model (PDAM). Under the PDAM model, Turtle Trees only incur a cost of $O\left(\log_{\rho}{\frac{N}{L}}\right)$ parallel I/Os per point query: one per node on the key’s root-to-leaf path. This makes point queries on Turtle Trees on-par with (non-buffered) B-trees.
Like other B-tree variants, range queries on Turtle Trees cost a point query (to find the starting key) plus $O\left(\frac{k}{B}\right)$ IOPs. There is also an in-memory, per-key cost of $\log_{2}{(\log_{2}{(N/L)})}$ to merge sorted runs, which is the same as a level-tiered LSM-tree. This per-key cost expression is derived via the use of a binary min-heap to select which of the $\log_{2}{N/L}$ levels of the tree contains the least-ordered key.
How $\chi$ effects cost analysis. We claim that, in general, the effect of $\chi$ is that keys skip the first $\log_{2}{\chi}$ buffer levels of the tree, which reduces the per-key write cost to:
@@ 3.4. Cost Analysis & Limitations
Limitations. The key limitation of TurtleTrees and $\chi$ is memory usage to support $\chi$ instead of read caching. But, $\chi$ can be tuned dynamically during runtime so $\chi$ need only be increased when needed. In return, TurtleTrees and $\chi$ pay less read-filter cost for the lifetime of data.
4. From TurtleTrees to TurtleKV
Figure 7 shows TurtleKVs high-level architecture. It is similar to systems based on log-structured B-trees and on LSM-trees (Figure 7). Per-key updates are inserted into an in-memory index, or MemTable, and appended to a write-ahead log (WAL) for durability. The purpose of the MemTable, as in LSM-tree designs, is to buffer updates until we have reached a minimum batch size. Once this happens, the checkpoint pipeline sorts and de-duplicate key/value pairs in the MemTable, producing a batch which is applied to the most recent checkpoint (which is a TurtleTree). This process is repeated until $\chi$ batches (MemTables) have been applied, at which point the modified portions of the TurtleTreeare serialized and written to storage. Afterward, the WAL is trimmed.
Queries are processed by first consulting the MemTable, which represents recent updates (up to some size threshold). If the key is found here, the query finishes. Otherwise, the query proceeds to check recent MemTables in most-recent order, one per batch applied to the in-memory pending checkpoint. If the key is not found in any of these, the most recent durable checkpoint is searched.
Implementation & parameters We implemented TurtleKV in approximately 30k lines of C++ code. TurtleKV’s static parameters include $B$ (the block size), $M$ the total memory budget, the filter bit rate ($FBR$), and the number of threads involved in various stages of updating Memtables and the checkpoint pipeline ($T$). In this work, we investigate the value of dynamically changing checkpoint distance ($\chi$) on demand, This parameter varies the delay between creating pages or batches and writing them to disk.
TurtleKV needs to address two high-level challenges. First, it must guarantee high throughput and low latency. Second, for $\chi$ to have value, it must guarantee that I/O, not memory or CPU is the bottleneck. We next describe TurtleKV’s key components in more detail and how we address these challenges for them.
Figure 7. High-level Architecture of TurtleKV
4.1. TurtleKV Component Details
Write-Ahead Log. TurtleKV writes all key updates to a shared change log file initially, to provide data durability in the event of program termination before the next checkpoint update. This file is divided into fixed-sized blocks, which internally use a slotted page format to store variable-sized per-key edits.
To avoid low performance due to contention when logging updates, we implement a separation-of-concerns strategy where each user thread writes key updates to a thread-local block buffer to minimize contention. A single background thread is responsible for flushing data to storage. The background flusher thread uses a lock-free, periodic polling scheme to steal ownership of buffers from foreground threads. The polling period can be adjusted to trade off between minimizing the latency for updates to become durable, and the throughput achieved by update batching.
Change log blocks may be configured large enough (e.g., ¿ 4kib) so that writing to storage is not atomic. Therefore the block header includes a hash of its contents. At recovery time, any block whose hash does not match is assumed to be partially written, and is ignored. Each update to the change log is assigned a unique, monotonic sequence number using an atomic counter. This can be used during recovery to maintain a consistent order of events across restarts.
MemTables. We use Adaptive Radix Trees (ARTs) as the core data structure for indexing buffered updates (Leis et al., 2013). This allows TurtleKV to maintain an ordered index for range queries, providing low latency. A concurrent hash table index can also be enabled at runtime to further speed point queries, at the cost of slightly higher insertions since the ART is still used.
To mitigate contention between multiple writers, TurtleKV uses optimistic lock coupling (OLC) (Leis et al., 2016). We also exploit the fact that MemTables do not need to support true key removals, since deletions must be merged into the checkpoint tree. This allows us to use a simple arena-based memory allocation scheme, with each thread allocating memory from the system in 64kib chunks which are then divided as needed into ART nodes. TurtleKV MemTables rely on WAL block buffers to store actual key/value data, further simplifying the job of the ART index, since it does not need to copy update data.
Checkpoint Pipeline. We use the parallelism of the underlying hardware to perform the needed work. A foreground thread first places finalized Memtables in a queue, which is consumed by a background thread. This thread sorts and compacts key/value pairs to produce new batches of updates. Each batch is placed in another queue which is consumed by the thread responsible for updating the checkpoint tree.
The checkpoint pipeline must also manage copying of data as sorted key/value pairs are merged/compacted within buffer levels and flushed down the tree. By increasing $\chi$, we expect incoming key updates to be pushed further down the tree before they are written to storage. If we naïvely copy key/value data into page buffers during this process at every step, most of this work will be wasted in building pages which are never written.
To address this challenge, Checkpoint tree updates defer the copying of key/value data until the last batch has been applied (according to the $\chi$ parameter). Merge/compactions on in-memory checkpoint tree update buffer levels and leaves therefore operate on pointers, to avoid data copies. The individual merge/compact operations are parallelized using a background thread pool. This allows write-heavy workloads that execute on only a few threads to take advantage of available CPU cores when possible.
When the final batch for a checkpoint has been applied, all new pages (leaf, node, and filter) are built concurrently using the same thread pool used for merging. Once all pages have been built, the checkpoint is handed via a third queue to the final thread, which writes the page data in parallel to storage. Page lifetime is controlled via reference counting. The reference counts, as well as root references to subsequent checkpoint trees, are appended to a write-ahead log using a two-phase commit protocol to ensure crash consistency.
Once a new checkpoint has been made durable, the checkpoint field of the KVStore is updated and the finalized MemTables representing the batches used to create the new checkpoint are pruned from the Deltas stack. Pruning the MemTables from Deltas allows the change log to be trimmed, releasing storage resources and ensuring that recovery times do not grow unbounded.
Page Cache. TurtleKV uses LLFS (Astolfi et al., 2024), MathWorks’ Low-Level File System library. This means that storage resources are organized into different pools (extents) of storage, one per page size. [^7] Typically one for nodes (4kb), one for leaves/segments (2mb-16mb), and one for filters (16kb-256kb). Each of these pools is represented in LLFS as a PageDevice.
While it is advantageous to use large (>2>2mb) page sizes for leaves when writing new checkpoint pages, reading entire leaf pages while processing a point query can introduce a severe latency penalty. To address this issue, we add a kind of virtual PageDevice to LLFS, called a sharded view. Sharded views assign page identifiers to slices of larger pages; e.g., a 4kb sharded view of a 16mb page device allows the loading/caching of 4kb aligned slices. This feature allows the point query code path to load and cache only the parts of a leaf page that are needed to answer a query.
Using sharded views allows the TurtleKV’s large leaf page design to perform competitively in low cache configurations when compared to the alternate Bϵ+-tree design in SplinterDB, which uses 24mb B+-trees composed of 4kb pages instead of a single large page. TurtleKV’s sharded page view design has the additional advantage of being able to answer key presence queries (return true or false depending on whether the key is present, instead of returning its value) more efficiently, since key and value data are stored in separate parts of a leaf.
Caching policy. Even if we optimize all the other steps so that write I/O is the bottleneck, we still need to achieve a high cache hit-rate when updating the TurtleTree checkpoint. This is because the I/O improvements we obtain from increasing $\chi$ may be offset by increased I/O from reading pages containing data to merge, compact, or flush. To address this issue, we use use a modified CLOCK (Jiang et al., 2005) cache replacement policy where each page has a priority value (instead of a single bit) which is decremented when the clock hand passes it. This allows TurtleKV to provide information to the lower layer about which pages to make an extra effort to keep in memory.
@@ 5. Evaluation
We claim that TurtleKV demonstrates how a key-value storage engine can be built to support low cost re-tuning at runtime, with a dynamic trade-off range that provides excellent resource utilization and competitive performance for write-heavy, read-heavy and mixed workloads. The goal of the experiments in this section is to support this claim.
We aim to answer the following questions:
  • Question 1: Does TurtleKV offer competitive throughput, write amplification, and space amplification for write-heavy workloads?
  • Question 2: Does TurtleKV offer superior latency and throughput in read-heavy workloads, dominated by point and range queries?
  • Question 3: What are the properties, both in cost and benefit, of runtime adjustments to checkpoint distance to select an optimal tuning point in response to workload?
Methodology. To answer these questions, we use the Yahoo Cloud Serving Benchmark (YCSB), comparing TurtleKV to other state-of-the-art key-value engines, each representing a different optimization target: RocksDB, a write-optimized and widely used LSM-tree-based system; and SplinterDB, a state-of-the-art hybrid-optimized key-value store. Splinter and Rocks are run mostly in their default configurations; direct I/O was enabled in RockDB in order to control the size of the block cache (this resulting in a slight improvement in overall throughput).
In our benchmarking, we use two record sizes, 128 bytes and 1024 bytes, 8 bytes of which is the key. Each workload is preceded by an initial load phase, where we populate the database with 48GiB of data, 400M 128-byte records or 50M 1024-byte records. We test both out-of-cache and in-cache workloads; for out-of-cache, we limit cache size to 16GiB, $1/3$ of the pre-loaded data size, and for in-cache we allow up to 48GiB of cached pages. In our parameter selection process, we found that for out-of-cache workloads, a relatively small number of threads is needed to reach the point where I/O becomes the bottleneck. Correspondingly, for in-cache, the bottleneck shifts to CPU and memory bandwidth. Consequently, our out-of-cache tests use 4 client workload threads, and in-cache tests use 16 threads. We also include results for single-threaded tests to show latency of sequentially issued requests.
@@
Hardware setup. Experiments were run on an AMD Ryzen Threadripper 7970x workstation (32 cores/64 threads), with 128 GB of DDR5-6000 registered ECC memory. The SSD used was an Intel P4800x 1.5TB, with 512 byte LBA configured, formatted with XFS. The P4800x is a data center grade Optane device supporting PCIe 3.0, rated for 2.4 GB/s read bandwidth and 2 GB/s write (550k/500k IOPs respectively). The system uses Ubuntu Linux 22.04 with kernel version 6.3. The CPU, while physically packaged as a single-socket part, was configured in the firmware (ASRock TRX50WS) to report its topology as a 4-node NUMA processor, one node per chiplet, each with 8 cores/16 threads.
(a) 400M $\times$ 128B, 1 thread
(b) 400M $\times$ 128B, 4 threads
(c) 400M $\times$ 128B, 16 threads
(d) 50M $\times$ 1024B, 1 thread
(e) 50M $\times$ 1024B, 4 threads
(f) 50M $\times$ 1024B, 16 threads
Figure 8. YCSB Results (higher is better): TurtleKV matches or exceeds performance of other engines on small records, and is competitive on large records. For 1024B records, TurtleKV performs better in higher cache configs due to separation of key and value data in large (16mib) leaf pages degrading query performance under cache pressure.
(a) 400M $\times$ 128B, Write Amp
(b) 50M $\times$ 1024B, Write Amp
(c) 400M $\times$ 128B, Space Amp
(d) 50M $\times$ 1024B, Space Amp
Figure 9. Write & Space Amplification (lower is better): TurtleKV has slightly higher WAF than SplinterDB, but uses less than half the space on disk.
(a) 400M $\times$ 128B, Workload E (scans)
(b) 50M $\times$ 1024B, Workload E (scans)
Figure 10. TurtleKV’s balanced compaction policy allows it to achieve high throughput when performing short range queries (<<100 records).

### v2

@@ 3.3.3. Write-Memory Trade-off Efficiency
We claim that, in general, the effect of $\chi$ is that keys skip the first $\log_{2}{\chi}$ buffer levels of the tree, which reduces the per-key write cost to:
@@
The key limitation of TurtleTree’s $\chi$ tuning knob, and of RM+WM tuning schemes in general, is that unlike I/O operations, memory is generally a bounded quantity. This means higher $\chi$ value will spend memory that could have been used for page and filter caching. Thus the zero-sum of fixed-size memory resources is effectively turned into a read-vs-write (RW) trade-off. However, unlike RW trade-offs that affect the structure of stored data, our scheme is flexible and forgiving, since it can be changed at runtime.
4. TurtleKV System Design
In this section we first give an overview of the system architecture and update/query paths in TurtleKV, then dive deeper into the design of two of the key ideas it implements: checkpoint distance tuning with Big MemTables, and parallel (multi-core) acceleration for TurtleTree (checkpoint) updates.
4.1. Basic Operation
TurtleKV is based on a conventional LSM-tree system architecture employing three high-level components: a write-ahead log (WAL) to ensure durability of incoming updates, an in-memory index and buffer of recent data (MemTable), and an on-disk checkpoint for long-term storage (TurtleTree). It is implemented in  32k lines of C++ code, and is available as open-sourced software (HttpsGithubcomMathworks2025).
4.1.1. Update Path
Incoming key/value data are copied to an internal memory buffer which is continuously flushed in the background to the WAL. Each update is assigned a unique sequence number so that the database state can be consistently reconstructed during recovery. After the key and value bytes are copied, they are inserted into an in-memory index (the active MemTable) on the user thread performing the update; as soon as this completes, the update is visible to all threads in the process.
Each MemTable is limited to a maximum size $\overline{M_{w}}$. When a thread attempts to perform a MemTable insertion, but fails because this limit would be exceeded, the active MemTable is finalized, or made read-only, and a new active MemTable is created to accept future updates. The finalized MemTable is handed off to a background thread, which splits it into 1 or more leaf-page-sized batches, each of which is applied to the current checkpoint TurtleTree. The user thread must wait if the background thread is already busy processing a previous finalized MemTable. After the hand-off succeeds, it will also wait until the total number of finalized MemTables in the system is at most 2. When the background thread finishes applying all batches for a MemTable to the page cache, it hands the list of new pages to a second background thread (the writer thread), which writes the new checkpoint data to disk. Once the new TurtleTree pages are durable, the writer thread swaps the finalized MemTable with a reference to the new checkpoint which subsumes it, possibly unblocking any application threads that were waiting for the finalized MemTable count to decrease.
This multi-stage update pipeline allows high system resource utilization by allowing the different stages (MemTable insertion, TurtleTree updates, and checkpoint page flushing) to be performed at the same time. By keeping the amount of queued data in the pipeline to a minimum, it also provides back-pressure so that no part of the system outpaces the others, in order to avoid unbounded tail latencies.
4.1.2. Query Path
Point queries are answered by first consulting the active MemTable. If the target key is found there, the query terminates and the answer is returned to the caller. Otherwise, the $\leq$2 finalized MemTables are consulted in newest-to-oldest order. If the key is not found there either, then the most recent checkpoint is queried as described in $\S$3.2.2.
TurtleKV supports either Bloom or Quotient filters (selected at compile-time) to optimize point queries. Each leaf page, whether at the bottom of the TurtleTree or referenced from a node buffer, has an associated filter page. Whenever the query algorithm demands that a leaf be searched for a given key, the associated filter is first queried, and only if it gives a positive response is the leaf loaded/checked.
When a leaf is queried, we first attempt to pin the whole leaf page in the cache; this will fail if the page isn’t in memory, without performing any I/O. If the whole leaf is not cached, then we access the first  64kb-sized slice of the page, which contains the header and trie index. From this information, we are able to determine which 4kb slice will contain the target key, if present.
4.2. Parallelization of TurtleTree Updates
4.2.1. The necessity of parallelized updates
TurtleTree checkpoint update is the most expensive operation in the TurtleKV update path by a substantial margin. Therefore it is the primary target of our optimization efforts to avoid in-memory bottlenecks, which is one of the 3 high-level requirements for TurtleKV ($\S$2, Table 1).
[…]
The challenge of TurtleTree updates is that they represent a potentially logarithmic amplification of work as compared to earlier pipeline stages, due to the $\log(N/L)$ levels through which update records must be merged/compacted to reach the leaf level. There is simply no way to keep up with load generated by earlier stages of the update path if we do not employ some kind of in-memory parallelization strategy.
4.2.2. Task-parallel vs Data-parallel
There are two basic approaches to this: allow concurrent updates to take place in different locations within the same tree, or somehow parallelize the work of sequentially applied batch updates. The former approach (task-parallel) is used by SplinterDB; TurtleKV takes the latter approach (data-parallel). The reason is twofold: first, there are many complications involved in implementing concurrent updates to Bϵ+-trees, and these are a superset of the complexities one faces when implementing concurrent B+-trees, a non-trivial task because of the competing concerns of root-to-leaf traversal for insertion, and leaf-to-root split/merge for re-balancing. The second reason is, as we explain below, that adopting a data-parallel approach allows us to better control load balancing across processors, leading to more consistent update performance with less variability.
4.2.3. Details of parallel updates
We observe that the most CPU-intensive operations in TurtleTree batch update are the key comparisons required to merge/compact level segments when a node receives an incoming batch from its parent and flushes an outgoing batch to a subtree. Merging multiple ordered sequences is easily done in parallel using multi-selection (deoOptimalParallelAlgorithm1994) on each sequence to determine the input and output ranges for different processors. By implementing this algorithm, we are able to scale the CPU work to multiple processors while ensuring the correctness of our implementation through simplicity of the code.
We also observe that the most memory-bandwidth expensive operation during checkpoint update is copying key-value data while serializing new compacted pages prior to writing them to storage. Our approach to optimizing this is reduce the total amount of data that must be copied by deferring all serialization until right before the checkpoint is finalized, and to serialize all leaf pages (and their corresponding filters) in parallel on different threads.
4.3. Checkpoint Distance & Big MemTables
4.3.1. Preliminaries: MemTable Implementation
The requirements for the MemTable data structure are scalability to high numbers of concurrent threads, low space overhead, ability to perform key-order scans, and CPU cache friendliness. There are several possible data structures which meet these criteria; for example, Wang et al. (wangBuildingBwTreeTakes2018) compare Bw-Tree, OpenBW-Tree, SkipList, Masstree, B+-tree, and Adaptive Radix Tree (ART) implementations for in-memory, multi-threaded performance and efficiency. TurtleKV implements ARTs for this purpose, because they outperform all other data structures in this comparison, and offer low memory usage and high CPU-cache hit rate across a variety of workloads. However, TurtleKV is careful to abstract this implementation choice so that other data structures can be swapped in if desired. For comparison, RocksDB uses concurrent SkipLists as the default MemTable, WiredTiger uses a combination of B+-tree and SkipLists for in-memory indexing, and SplinterDB uses concurrent B+-trees as its MemTable and branches. The TurtleKV ART implementation is roughly the same size (lines of code) as RocksDB’s SkipList.
(a) KOps/Second, 16gb cache
(b) KOps/Second, 16gb cache
(c) KOps/Second, 48gb cache
(d) KOps/Second, 48gb cache
[…]
4.3.2. Checkpoint Distance: A Design Dilemma
TurtleKV’s write-memory (WM) tuning knob is checkpoint distance; this is the total size of the updates between two checkpoints. This can be varied at runtime to target different levels of write-optimization at the expense of higher memory footprint devoted to updates. Recall from our case study ($\S$2.2.2), that we find there are two knobs offered by RocksDB to trade write amplification directly against memory: MemTable size $\overline{M_{w}}$ and MemTable count $|M_{w}|$. The theoretical reason behind their effectiveness is that they force the first $O(\log_{F}(M))$ compactions for a given key $x$ to happen in memory, before that key is ever written to an SSTable. Varying either parameter while holding the other constant is equally effective, since $M=\overline{M_{w}}\cdot|M_{w}|$. The reason to prefer increasing $\overline{M_{w}}$ to $|M_{w}|$ as a tuning knob is that MemTables’ contribution to query cost in a RocksDB-like design is $O(|M_{w}|\cdot\log(\overline{M_{w}}))$, assuming a MemTable data structure that supports optimal logarithmic query cost.
If we apply this same logic to TurtleKV’s design, however, we encounter a dilemma. We want the on-disk structure of checkpoint TurtleTrees to be independent of checkpoint distance, so that we get a true WM tuning knob; this is possible only if we keep $\overline{M_{w}}$ constant and vary the number of batches $|M_{w}|$ in between checkpoints. Thus we have a tug-of-war between the concerns of in-memory (MemTable) queries on newer data, and queries against checkpoints (TurtleTrees) in the face of varied past tuning point selection.
4.3.3. The Solution: Big MemTables
We resolve this tug-of-war by setting the size limit of a MemTable when it is created to the current checkpoint distance, and breaking up finalized MemTables into many leaf-page sized batches in order to update the TurtleTree. We do this via a least-to-greatest key-ordered scan of the MemTable. This design is called Big MemTables in TurtleKV.
Big MemTables help to optimize TurtleKV’s performance in multiple ways. First, by scaling up the size of TurtleKV’s in-memory index on recent updates, we speed up queries on ”hot keys.” Second, by allowing a larger set of updates to accumulate in memory before doing a scan to break the MemTable into batches, we increase the skew in key distribution when applying batches to the checkpoint. TurtleTree batch update ($\S$3.2.1), which flushes batches down the tree as soon as the minimum size threshold for a subtree is reached (not when the buffer fills up), transparently leverages this key skew to move data down the tree faster, so long as doing so will not hurt efficiency (i.e., so long as we preserve the property that merges are always performed on equal-sized inputs). We note that SplinterDB’s flush-then-compact policy achieves a similar effect, though it has the downside of increased space amplification due to the fact that a flushed branch may end up referenced in many different subtrees despite only containing a small amount of live data for each reference.
Finally, Big MemTables further motivate our selection of ARTs as MemTable data structure. Because the structure of radix trees allow for segments of indexed keys to be shared, it is possible for the per-key overhead of ARTs to fall below the average key length (leisAdaptiveRadixTree2013); this efficiency is often increased as the total set of keys in an ART scales. Thus by increasing $\overline{M_{w}}$ and then splitting buffered updates into key-skewed batches, we not only play to the strengths of TurtleTrees, but to those of ARTs as well.
@@ 5. Evaluation
In this section, we explore the end-to-end performance of TurtleKV, comparing it to WiredTiger, RocksDB, and SplinterDB. We provide data to support the claims we make about TurtleKV and its novel data structure TurtleTrees, and our claims about the necessity of write-memory (WM) trade-off tuning more generally.
(a) KOps/Second, 16gb cache
(b) KOps/Second, 16gb cache
(c) KOps/Second, 48gb cache
(d) KOps/Second, 48gb cache
(e) WAF vs Checkpoint Dist. (MB)
Figure 9. TurtleKV WM Tuning Sensitivity
5.1. Experimental Setup
5.1.1. Hardware setup
@@
Experiments were run on an AMD Ryzen Threadripper 7970x workstation (32 cores/64 threads), with 128 GB of DDR5-6000 registered ECC memory. The SSD used was an Intel P4800x 1.5TB, with 512 byte LBA configured, formatted with XFS. The P4800x is a data center grade Optane device supporting PCIe 3.0, rated for 2.4 GB/s read bandwidth and 2 GB/s write (550k/500k IOPs respectively). The system uses Ubuntu Linux 22.04 with kernel version 6.3. The CPU, while physically packaged as a single-socket part, was configured in the firmware (ASRock TRX50WS) to report its topology as a 4-node NUMA processor, one node per chiplet, each with 8 cores/16 threads.
5.1.2. Software/Benchmark Configuration
We use the Yahoo Cloud Serving Benchmark (YCSB)(cooperBenchmarkingCloudServing2010), comparing TurtleKV to other state-of-the-art key-value engines, each representing a different optimization target: WiredTiger, a read-optimized B+-tree-based engine; RocksDB, a write-optimized and widely used LSM-tree-based system; and SplinterDB, a state-of-the-art hybrid-optimized key-value store. Each YCSB run inserts 400M records comprised of 8 byte keys and 120 byte values (Load), then performs 150M operations comprised of 50% updates/50% gets (workload A), another 150M operations comprised of 5% updates/95% gets (workload B), 150M gets (workload C), 15M operations comprised of 95% scans (len=1..100)/5% updates (workload E), and then a final 150M operations comprised of 50% read-modify-writes/50% gets (workload F). The total size of the initial data set is 48gb; to evaluate both out-of-cache and in-cache performance, we run the benchmark twice for each database, once with cache size 16gb ($\frac{1}{3}$ of data size) and again with 48gb cache (large enough to fit the raw data).
5.1.3. Database Configurations
The configuration used for each database is informed by the case studies described in $\S$2.2. For WiredTiger, we give a generous allocation to write buffering in cache, setting eviction dirty target to half the cache size and eviction dirty trigger to 95% of the cache. We turn on transaction-level logging for durability, and disable automatic checkpointing for our tests. For RocksDB, we set parallelism equal to the number of cores on the test machine (64); we also set the maximum number of parallel sub-compactions to this number. We set the write buffer size equal to 64mb, since in our testing we did not observe an end-to-end advantage to larger values (despite lower write amplification). We enable the block cache explicitly in order to accurately track the memory RocksDB uses for page caching (otherwise it uses the kernel file cache), and also enable direct I/O, which we observe to be slightly faster. SplinterDB is configured with mostly default values. We set the maximum key size to 8 bytes to match our test workload, and configure the number of background threads to 2 for memtable tasks, 24 for normal tasks, following the guidance of the SplinterDB tuning guide.
(a) Insert, (16gb)
(b) Update (16gb)
(c) Get (16gb)
(d) Scan (16gb)
(e) Insert (48gb)
(f) Update (48gb)
(g) Get (48gb)
(h) Scan (48gb)
Figure 10. YCSB, Per-Operation Latency Distributions for 16gb and 48gb Cache Sizes